OpenStudy (anonymous):

1a) Determine exact radian measure for each angle, given that 45 degrees is exactly pie/4 rad (answer 5pie/4) 2a)Determine exact radian measure for each angle, given that 45 degrees is exactly pie/4 rad (answer pie/8)

OpenStudy (anonymous):

Kind of confused here

OpenStudy (anonymous):

First, is there a picture that goes with this? Second, have you studied radians before?

OpenStudy (anonymous):

No pictures, and yes I have but not very good at it

OpenStudy (anonymous):

In a circle, there are 360º and there are \[2\pi\] radians; so for this problem you can set up the conversion \[ A \ \textrm{radians} \Big( \frac{360º}{2\pi \ \textrm{radians}} \Big) = Bº\]

OpenStudy (anonymous):

Well, there is one problem - they don't give you the num bers they want you to convert....

OpenStudy (anonymous):

that's kind of an issue :/

OpenStudy (anonymous):

oh sorry for 1a) it's 22.5 degrees

OpenStudy (anonymous):

for 2a) it's 22.5 aswell

OpenStudy (anonymous):

sure they do, they are giving you the angle in radians,